In a student laboratory exercise, the wavelength of a 36,300 Hz ultrasound wave is measured to be 0.970 cm. Find the air temperature (in K).
1
Expert's answer
2019-04-30T09:52:08-0400

The speed of sound in the laboratory can be calculated as


"v=l\/t=l\u03d1"

 

Where 

"l=0.970 cm=0.0097 m"

and

"\u03d1= 36300 Hz"

Empirically the dependence of speed of sound in the air from the air temperature had been figured out to be approximated as


"v (m\/s) = 331.5 + 0.60 T(\u00b0C)"

Where T is in degrees of Celsius

We can then do the following

"331.5 + 0.60 T=l\u03d1"

"T=((l\u03d1-331.5))\/0.6"

Finally the air temperature in Kelvin can be calculated as 


"T_K=T+273.3=((l\u03d1-331.5))\/0.6+273.3=307.65 K"